# arXiv:2402.18042v3[math.AG] 01 Mar 2026

### v3

@@ 3 A relation between the Gersten-type conjecture and motivic cohomology groups
Let $U$ be the generic fiber of $\operatorname{Spec}(R)$ and $N$ an integer larger than or equal to $n$. Then we have an isomorphism
@@ 4.1 Equi-characteristic cases
Let $A$ be a local ring $\mathcal{O}_{Y,y}$ of $Y$ at a point $y\in Y$. It suffices to prove the quasi-isomorphism (67) for $\operatorname{Spec}(A)$. We prove the statement by induction on $\operatorname{dim}(A)$.
Assume that $\operatorname{dim}(A)=0,$ that is, $A$ is a field. Then we have the quasi-isomorphism (67) by [7, p.774, Theorem 1.2.2] and [30].
@@
Let $j:U\to\operatorname{Spec}(A)$ be an open immersion of $\operatorname{Spec}(A)$ such that $U$ is a regular scheme. Let $\iota:Z\to\operatorname{Spec}(A)$ be a closed subscheme of codimension $c\geq 1$ with open complement $j:U\to\operatorname{Spec}(A)$. Since $A$ is an excellent ring, we are able to choose such a non-empty open subset $U$ of $\operatorname{Spec}(A)$.
@@
where the maps ()(*) and ()(**) are quasi-isomorphisms by the quasi-isomorphism (68) and [7, p.776, Proposition 2.2. a)]. Since the homomorphism
@@
Let $A$ be a local ring of a normal crossing variety over the spectrum of a field $k$. Let $n\geq 0$, $l>0$ be integers with $(l,\operatorname{char}(k))=1$. Then we have isomorphisms
@@
for $x\in\operatorname{Spec}(A)$ and any integer $q$ where $A_{x}$ is the local ring of $A$ at a point $x\in\operatorname{Spec}(A)$.
@@
Finally, we prove the isomorphism (73). By [24, Proposition 2.1], it suffices to prove the isomorphism (73) in the case where $((\operatorname{Spec}(A))^{(0)})<n+1$. Let $A_{\bar{x}}$ be the strict henselian of $A$ at a point $x\in\operatorname{Spec}(A)$. If $l^{\prime}$ is a prime number with $(l^{\prime},\operatorname{char}(k))=1$, we have $\operatorname{cd}_{l^{\prime}}(A_{\bar{x}})=0$. So we have isomorphisms
@@ 4.2 Mixed characteristic cases
Let $\mathfrak{X}$ be a semistable family over the spectrum of a discrete valuation ring $A$ of mixed-characteristic $(0,p)$. Then we have quasi-isomorphisms
@@
and $l$ be a positive integer which is invertible in $A$. Suppose that $A$ contains $l$-th roots of unity. Then there exists an exact sequence
@@ 4.3 Application
Let $B$ be a henselian valuation ring of mixed characeristic $(0,p)$, $\pi$ a prime element of $B$ and $C=B[T_{0},\cdots,T_{N}]/(T_{0}\cdots T_{a}-\pi)$ for $a\leq N$. Then we observe a relation between $\Gamma\left(C,R^{n+1}\alpha_{*}\mu_{l}^{\otimes n}\right)$ and $\Gamma\left(C/(\pi),R^{n+1}\alpha_{*}\mu_{l}^{\otimes n}\right)$ for an integer $l$ which is prime to $p$.
@@
for $0\leq a\leq N$ and $i:\operatorname{Spec}(C/(\pi))\to\operatorname{Spec}(C)$ be the closed fiber of $\operatorname{Spec}(C)$ and $l$ an integer which is prime to $p$. Then we have the composition
@@
by [7, p.774, Theorem 1.2.4] and the absolute purity conjecture ([5]) where the sequences are exact. Since $A$ is a henselian discrete valuation ring, the homomorphism
@@
Let $B$ be a henselian discrete valuation ring of mixed characteristic $(0,p)$ and $\pi$ a prime element of $B$. Let $n$ be a non-negative integer and $l$ an integer which is prime to $p$. Let
@@
Let $A$ be a henselian local ring with $\operatorname{dim}(A)=1$ and $l$ an integer which is prime to $\operatorname{char}(A)$. Suppose that the maximal ideal of $A$ is principal. Then the homomorphism
@@ 5.1 A generalization of Artin’s theorem
Let the notations be the same as in Theorem 5.5.
@@
    Let $l$ be an integer which is prime to $p$. Then the natural map
@@
    for integers $s\geq 0$ and $t\geq 2$ by the same argument as in the proof of Theorem 5.5. Since we have quasi-isomorphisms

### v4

@@ 3 A relation between the Gersten-type conjecture and motivic cohomology groups
Let $U$ be the generic fiber of $\operatorname{Spec}(R)$ and NN an integer larger than or equal to $n$. Then we have an isomorphism
@@ 4.1 Equi-characteristic cases
Let AA be a local ring $\mathcal{O}_{Y,y}$ of $Y$ at a point $y\in Y$. It suffices to prove the quasi-isomorphism (67) for $\operatorname{Spec}(A)$. We prove the statement by induction on $\operatorname{dim}(A)$.
Assume that $\operatorname{dim}(A)=0,$ that is, AA is a field. Then we have the quasi-isomorphism (67) by [7, p.774, Theorem 1.2.2] and [30].
@@
Let $j:U\to\operatorname{Spec}(A)$ be an open immersion of $\operatorname{Spec}(A)$ such that $U$ is a regular scheme. Let $\iota:Z\to\operatorname{Spec}(A)$ be a closed subscheme of codimension $c\geq 1$ with open complement $j:U\to\operatorname{Spec}(A)$. Since AA is an excellent ring, we are able to choose such a non-empty open subset $U$ of $\operatorname{Spec}(A)$.
@@
where the maps ()(*) and ()(**) are quasi-isomorphisms by the quasi-isomorphism (68) and [7, p.776, Proposition $2.$ a)]. Since the homomorphism
@@
Let AA be a local ring of a normal crossing variety over the spectrum of a field $k$. Let $n\geq 0$, $l>0$ be integers with $(l,\operatorname{char}(k))=1$. Then we have isomorphisms
@@
for $x\in\operatorname{Spec}(A)$ and any integer $q$ where $A_{x}$ is the local ring of AA at a point $x\in\operatorname{Spec}(A)$.
@@
Finally, we prove the isomorphism (73). By [24, Proposition 2.1], it suffices to prove the isomorphism (73) in the case where $((\operatorname{Spec}(A))^{(0)})<n+1$. Let $A_{\bar{x}}$ be the strict henselian of AA at a point $x\in\operatorname{Spec}(A)$. If $l^{\prime}$ is a prime number with $(l^{\prime},\operatorname{char}(k))=1$, we have $\operatorname{cd}_{l^{\prime}}(A_{\bar{x}})=0$. So we have isomorphisms
@@ 4.2 Mixed characteristic cases
Let $\mathfrak{X}$ be a semistable family over the spectrum of a discrete valuation ring AA of mixed-characteristic $(0,p)$. Then we have quasi-isomorphisms
@@
and $l$ be a positive integer which is invertible in AA. Suppose that AA contains $l$-th roots of unity. Then there exists an exact sequence
@@ 4.3 Application
Let $B$ be a henselian valuation ring of mixed characeristic $(0,p)$, $\pi$ a prime element of $B$ and $C=B[T_{0},\cdots,T_{N}]/(T_{0}\cdots T_{a}-\pi)$ for $a\leq N$. Then we observe a relation between $\Gamma\left(C,R^{n+1}\alpha_{*}\mu_{l}^{\otimes n}\right)$ and $\Gamma\left(C/(\pi),R^{n+1}\alpha_{*}\mu_{l}^{\otimes n}\right)$ for an integer $l$ which is prime to pp.
@@
for $0\leq a\leq N$ and $i:\operatorname{Spec}(C/(\pi))\to\operatorname{Spec}(C)$ be the closed fiber of $\operatorname{Spec}(C)$ and $l$ an integer which is prime to pp. Then we have the composition
@@
by [7, p.774, Theorem 1.2.4] and the absolute purity conjecture ([5]) where the sequences are exact. Since AA is a henselian discrete valuation ring, the homomorphism
@@
Let $B$ be a henselian discrete valuation ring of mixed characteristic $(0,p)$ and $\pi$ a prime element of $B$. Let $n$ be a non-negative integer and $l$ an integer which is prime to pp. Let
@@
Let AA be a henselian local ring with $\operatorname{dim}(A)=1$ and $l$ an integer which is prime to $\operatorname{char}(A)$. Suppose that the maximal ideal of AA is principal. Then the homomorphism
@@ 5.1 A generalization of Artin’s theorem
Let the notations be the same as in Theorem $5.$
@@
    Let $l$ be an integer which is prime to pp. Then the natural map
@@
    for integers $s\geq 0$ and $t\geq 2$ by the same argument as in the proof of Theorem $5.$ Since we have quasi-isomorphisms